OpenStudy (anonymous):

6 – 4x = 4(3 – x) – 6? I need help for

OpenStudy (anonymous):

6-4x=21-4x-6 6-4x=15-4x x=0 here, x must be 0 for left hand side to equal right hand side

OpenStudy (anonymous):

Expand of the right hand side as artofspeed suggests. Take care though, their expansion is not quite correct. once you have the right hand side expanded, you can isolate x and solve.